# GASLITEing the Retrieval: Exploring Vulnerabilities in Dense Embedding-based Search

### v1

@@ 4 Method
4.1 Formalizing the Adversary Objective
The adversary seeks to create textual adversarial passages, , maximizing the probability of retrieving at least one adversarial passage () ranked in the top- results, over queries () sampled from the targeted distribution . Formally:
@@ 4.1 Formalizing the Adversary Objective
To simplify the objective, we assume no constraints in and that =1, relaxing these assumptions later. Following our threat model, attackers cannot use and may employ available sample queries (). We can then estimate the objective as (see App. B for detailed derivation):
@@
This estimated objective (Eq. 2) maximizes the alignment between the controlled adversarial passage and the centroid of the targeted query distribution , suggesting that should reflect the “average” semantic of the targeted queries. Eq. 2 provides a compact, query-count-agnostic objective of a single-vector inversion, which we optimize efficiently with GASLITE (§4.2), and use to form an hypothetical baseline attack achieving this objective (perfect attack in §6.4).
@@
Constraining the Objective. To ensure informativeness (§3), we construct by concatenating a fixed prefix containing the malicious information () with an optimized trigger: . The attack optimizes the trigger while keeping the prefix intact. For additional stealth, we can further constrain the attack (e.g., require a fluent trigger) by incorporating relevant terms (e.g., text perplexity) into the objective (as we do in defenses evaluation; App. G).
@@
Generalizing for Larger Budgets. Finding the optimal solution for multi-budget settings is generally -Hard (reduction to Set Cover in App. D.2). Thus, to accommodate larger budgets (), we partition the query set into subsets and attack each separately, optimizing Eq. 2 per subset. (see Alg. 2 in App. D.1). After empirical evaluation of various partitioning methods, including different clustering algorithms and variants of greedy set cover approximation (Korte & Vygen, 2012) we found -means (Lloyd, 1982) to best use a given budget—although, in specific low-budget cases, the greedy algorithm performs slightly better (App. D.3). Hence, we choose -means as the query-partition method and further motivate it with a desired theoretical property it holds (App. D.3).
4.2 Optimization With GASLITE
@@
To systematically optimize toward the objective in Eq. 2, we introduce GASLITE (Gradient-based Approximated Search for maLIcious Text Embeddings), a multi-coordinate ascent gradient-based algorithm that iteratively refines textual triggers to maximize the similarity between adversarial passages and target queries within the embedding space (Eq. 2). We now describe GASLITE (Alg. 1), outlining the critical design decisions, and demonstrate its superiority as an optimizer for attacking retrieval compared to prior optimizers (Fig. 2), with more comprehensive evaluation to follow (§6).
While continuous optimization of Eq. 2 is straightforward, involving only the computation of the query centroid, the challenge lies in finding a text , in a discrete space, that satisfies this objective. Prior attacks on NLP models (§2) have addressed this by leveraging gradient information to reduce the search space. Specifically, Ebrahimi et al. (2018) proposed a well-established mathematical scheme, which we employ in our attack and is at the foundation of several LLM jailbreaks (GCG (Zou et al., 2023), ARCA (Jones et al., 2023)) and retrieval attacks (Cor.Pois. (Zhong et al., 2023)). This scheme: (1) computes the gradient of the objective w.r.t. the input to estimate a linear approximation (i.e., first-order Taylor) over all tokens; (2) uses this approximation to identify promising candidates for token substitutions (i.e., ones that are likely to increase the objective); and (3) evaluates the exact objective for these candidates (via forward passes), and performs the best substitution evaluated.
@@
GASLITE Algorithm. GASLITE improves upon past methods in three key ways, each empirically critical: (i) it refines the objective approximation by averaging linear approximations around multiple token-substitutions sampled from a vast vocabulary, extending ARCA; (ii) it substitutes multiple token in various positions per iteration, reducing the required backward passes per substitution; and (iii) it maintains the performance of adversarial passages throughout the attack (i.e., the text produced from decoding optimized input tokens), as attackers inserts adversarial text not tokens to the corpus (§3). Ablating each of these causes 5–16% drop in attack success (App. F.2), and comparison to prior optimizers show GASLITE’s superior speed and efficacy (Fig. 2; App. F.1).
@@
Alg. 1 outlines our method for generating an optimized text trigger of tokens, maximizing Eq. 2 for a retrieval model and query set . We start by calculating the target vector per Eq. 2 (L1) and initialize the trigger with arbitrary text (L2). Next, for iterations, we calculate the linear approximation of the objective, averaged on random single-token flips on (L4–5). Then, we randomly choose a subset of token positions (L6) where we perform the token substitutions, and, for each position, we use the linear approximation to identify promising token substitutions (L8-9), filter irreversible tokenizations (L10), and evaluate the exact objective on the remaining candidate triggers, picking the objective-maximizing token substitution to update (L11).
Input: embedding model, set of textual queries, trigger length , , , ,
5 Experimental Setup
We develop an extensive setup to test the susceptibility of popular, leading retrievers to SEO under varied assumptions and compare GASLITE to adequate baselines (see App. E for more details).
Models. We evaluate diverse embedding-based retrievers (Tab. E.1): MiniLM (Wang et al., 2020); E5 (Wang et al., 2022); Arctic (Merrick et al., 2024); Contriever and Contriever-MS (Izacard et al., 2021), ANCE (Xiong et al., 2020); GTR-T5 (Ni et al., 2022); and MPNet (Song et al., 2020b). We select these models based on performance (per retrieval benchmarks (Muennighoff et al., 2023)), popularity (per HuggingFace’s downloads and open-source usage ), diverse architectures (i.e., backbone model, pooling method, and similarity function), usage in prior work, and size (specifically with 110M parameters, the size of BERT-base), as efficiency is a desired property when working with large corpora. We include an additional evaluation with LLM-based embeddings in App. H.4.
Datasets. Focusing on retrieval for search, we use the MSMARCO passage retrieval dataset (Bajaj et al., 2016), containing a corpus of 8.8M passages and 0.5M real search queries, which we poison and target in attacks, respectively. For , we sample toxic statements from ToxiGen (Hartvigsen et al., 2022), and for concept-specific content, we use GPT4 (OpenAI, 2024) to create negative statements. We also validate results on the NQ dataset (Kwiatkowski et al., 2019) (App. H.3).
Our Attack. To simulate worst-case attacks while ensuring passages remain within benign passage length (App. F.3), we evaluate GASLITE for crafting passages where a malicious prefix is fixed, followed by a trigger of length (i.e., ). We extend GASLITE to a multi-budget attack using -means for query partitioning (§4.1). For additional hyperparamters we set , , and , as elaborated in App. F.3.
Baselines. We consider two naïve baselines and a major prior work for comparison, all performing an informative attack (). First, as a control, in Only we attack with the chosen alone (). Second, following a common SEO baseline (Zuze & Weideman, 2013; Zou et al., 2024; Shafran et al., 2024), we use stuffing—i.e., filling the with sample queries (App. E). Third, we employ Cor.Pois. attack (Zhong et al., 2023) as a strong baseline, using the original implementation, while allowing the attack to operate under its more permissive threat model where the attacker can access the corpus. For fair evaluation, all methods perform query partitioning using -means (§4.1) and share length ().
Metrics. As informativeness is inherent in the attacks ( serves as a prefix in crafted passages), we measure the attack success in terms of visibility. To this end, we adopt the well-established metric of appeared@k (Zhong et al., 2023; Song et al., 2020a), measuring the proportion of queries for which at least one adversarial passage () appears in the top- results; we typically set =10, per common search apps (e.g., the first page of Google search commonly displays 10 results), taking measurements over held-out queries (except in Knows All attacks in §6.1).
6 Evaluating The Susceptibility of Embedding-based Search
We evaluate our attack in three settings (§6.1–6.3), each corresponding to a different type and attacker knowledge of the targeted query distribution (per §3). We chiefly focus on the Knows What setting (§6.2), as it better reflects realistic SEO. Next, in §6.4, we assess the results using the hypothetical perfect attack. Appendices provide further results, including comparison across threat models (Tab. 6), a study on defenses and adapting GASLITE to them (including making adversarial suffixes fluent; App. G), transferability to black-box models (App. H.7), and generalizability to unseen datasets (App. H.5.2).
6.1 “Knows All”
@@ 6 Evaluating The Susceptibility of Embedding-based Search
Results. Tab. 1 shows that while the content alone ( Only) never appears in the top-10 results, simply appending the query (stuffing) boosts visibility to 30% avg. appeared@10. Cor.Pois. (Zhong et al., 2023) underperforms, sometimes worse than the naïve stuffing; we find it is mainly, albeit not only, due to generating adversarial tokens that, once decoded into text and tokenized to model input, result in vastly different tokens than the ones optimized (see App. I). Importantly, carefully designing the suffix with GASLITE renders it optimally visible, consistently ranked as the top-1 passage for each query; we attribute this to GASLITE’s passages achieving an exceptionally high vector similarity with the target query (Tab. 7)—for some dot-product models (e.g., Contriever) this similarity is twice that of baselines, a phenomenon we discuss later (§7).
6.2 “Knows What”
@@
Setup. We test attacks’ ability to increase visibility of specific (e.g., a negative Harry Potter review) across queries on a targeted concept (e.g., Potter). To test this setting, we choose eight recurring concepts of varying semantics and frequency, from MSMARCO (see App. E.4). Per our threat model (§3), we employ a sample (50%) of concept-related queries—albeit this can be relaxed by generating synthetic queries (App. H.5)—leaving a held out set (50%) of concept-related queries for evaluation with varying budget sizes (). Tab. 9 and Tab. 10 list examples of crafted passages.
@@
Results. Fig. 3 shows GASLITE outperforms all baselines, increasing appeared@10 by 40%. Unlike single-query attacks (§6.1), stuffing fails here (except with Contriever; §7) highlighting the increased attack difficulty compared to Knows All. Note that GASLITE remains superior even when evaluated under a more permissive, less realistic threat model, measuring the success directly on the crafted input tokens instead of text (Fig. 16, App. H.2). Fig. 4 demonstrates that attack success increases along the attack budget, with insertion of =10 passages to 8.8M-sized corpus (a poison rate of 0.0001%) sufficing for 50% avg. appeared@10 on 6/9 retrievers for unknown concept-related queries.
@@
Consistent with Zhong et al.’s (2023) findings, we find Contriever models highly susceptible (Figs. 3–4), with a single adversarial passage achieving 100% appeared@10. Other models show varying success (see §7), yet GASLITE’s results approach the optimal solution of Eq. 2 (see §6.4). Last, we observe GASLITE attains 100% appeared@100 on all models for (Fig. 18, App. H.2), indicating substantial content promotion, even if not into the top-10 results.
6.3 “Knows [Almost] Nothing”
@@
Setup. We test attacks’ potency when targeting general, unknown queries from a wide and diverse query distribution; a setting Zhong et al. (2023) studied. To this end, we randomly sample 5% of MSMARCO’s training queries (25K queries), made available for attacks with budgets . We evaluate on MSMARCO’s entire, diverse evaluation queries (7K), held out from the attack.
@@
Results. GASLITE markedly outperforms the baselines (Tab. 11): naïve methods ( only, stuffing) fail to achieve any visibility in top-10 results (or even in top-100) and Cor.Pois. achieves appeared@10 of 6% (Tab. 11). GASLITE’s performance significantly increases along the budget size, from 0% appeared@10 for a single-passage budget (Fig. 19, App. H.3) to appeared@10 of 5%-20% with =100, for most models (Fig. 5). The potential performance for a further increased budget size is explored in §6.4. Consistent with Knows What evaluation (§6.2), we observe variance in attack success across models (see §7).
6.4 The perfect Attack

### v2

@@ 4. Method
4.1. Formalizing the Adversary Objective
The adversary seeks to create textual adversarial passages, , maximizing the probability of retrieving at least one adversarial passage () ranked in the top- results, over queries () sampled from the targeted distribution . This corresponds to the measure we later refer to as appeared@ (§5). Formally, can be written as:
@@ 4.1. Formalizing the Adversary Objective
To simplify the objective, we assume no constraints in and that =1, relaxing these assumptions later. Following our threat model, attackers cannot use , and may employ available sample queries (). We can then estimate the objective as (see App. B for detailed derivation):
@@
This estimated objective (Eq. 2) maximizes the alignment between the controlled adversarial passage and the centroid of the targeted query distribution , suggesting that should reflect the “average” semantic of the targeted queries. Notably, fully maximizing the objective (Eq. 2) guarantees a visibility of the adversarial passage in the top- results (Eq. 1), and empirically, the higher the objective, the stronger the passage’s visibility (Fig. 12 in App. B). Eq. 2 provides a compact, query-count-agnostic, corpus-agnostic objective of a single-vector inversion, which we optimize efficiently with GASLITE (§4.2), and use to form a hypothetical baseline attack achieving this objective (perfect attack in §6.4).
@@
Constraining the Objective. To ensure informativeness (§3), we construct by concatenating a fixed prefix containing the malicious information () with an optimized trigger: . The attack optimizes the trigger while keeping the prefix intact. We place the malicious information as a prefix, exploiting the Primacy Effect, where early content more strongly influences readers (Paul and Weiss, 2013). For added stealth, we can further constrain the attack (e.g., require fluent triggers) by integrating relevant terms (e.g., text perplexity) into the objective (as we do in defenses evaluation; §7).
@@
Generalizing for Larger Budgets. Finding the optimal solution for multi-budget settings is generally -Hard (reduction to Set Cover in App. D.2), to accommodate larger budgets (), we partition the query set into subsets and attack each separately, optimizing Eq. 2 per subset (see Alg. 2 in App. D.1). After empirical evaluation of various partitioning methods (see App. D.3), including different clustering algorithms and variants of greedy set cover approximation (Korte and Vygen, 2012), we found -means (Lloyd, 1982), a method also used in past attacks (Zhong et al., 2023), to best use a given budget—although, in specific low-budget cases, the greedy algorithm performs slightly better (App. D.3). Hence, we choose -means as the query-partition method and further motivate it with a desired theoretical property it holds—maximizing the in-cluster pairwise similarity (App. D.3).
4.2. Optimization With GASLITE
@@
To systematically optimize toward the objective in Eq. 2, we introduce GASLITE (Gradient-based Approximated Search for maLIcious Text Embeddings), a multi-coordinate ascent gradient-based algorithm that iteratively refines textual triggers to maximize the similarity between adversarial passages and target queries within the embedding space (Eq. 2). We now describe GASLITE (Alg. 1), outlining the critical design decisions, and demonstrate its superiority as an optimizer for attacking retrieval compared to prior optimizers (Fig. 2; App. F), with more comprehensive evaluation to follow (§6).
Input: embedding model, set of textual queries, trigger length , , , ,
@@
While continuous optimization of Eq. 2 is straightforward, involving only the computation of the queries’ centroid, the challenge lies in finding a text , in a discrete space, that satisfies this objective. Prior attacks on NLP models (§2) have addressed this by leveraging gradient information to reduce the search space. Specifically, Ebrahimi et al. (Ebrahimi et al., 2018) proposed a well-established mathematical scheme (namely, HotFlip), which we employ in our attack and is at the foundation of several LLM jailbreaks (GCG (Zou et al., 2023), ARCA (Jones et al., 2023)) and retrieval attacks (Cor.Pois. (Zhong et al., 2023)). This scheme: (1) computes the gradient of the objective w.r.t. the input to estimate a linear approximation (i.e., first-order Taylor) over all tokens; (2) uses this approximation to identify promising candidates for token substitutions (i.e., ones that are likely to increase the objective); and (3) evaluates the exact objective for these candidates (via forward passes), and performs the best substitution evaluated.
@@
GASLITE Algorithm. GASLITE enhances this scheme in three key ways, each empirically critical for a successful retrieval attack: (i) it performs a greedy search for a multi-token substitution, per iteration, thus reducing the required backward passes per substitution (Labs, 2024); (ii) it refines the objective approximation by averaging linear approximations around multiple token-substitutions sampled from a vast vocabulary, extending ARCA (Jones et al., 2023); and (iii) it maintains the performance of adversarial passages throughout the attack (i.e., the text produced from decoding optimized input tokens), as attackers inserts adversarial text, not tokens, to the corpus (§3).555Detokenization is not an injective function—different token lists can be decoded to the same string (see App. H). Our attack ensures adversarial passages correspond to the optimized token lists (upon tokenization). Ablating each of these causes 5–16% drop in attack success (§F.2).
@@
Alg. 1 outlines our method for generating an optimized text trigger of tokens, maximizing Eq. 2 for a retrieval model and query set . We start by calculating the target vector per Eq. 2 (L1) and initialize the trigger with arbitrary text (L2). Next, for iterations, we calculate the linear approximation of the objective, averaged on random single-token flips on (L4–5). Then, we randomly choose a subset of token positions (L6), where we perform a greedy search over the token substitutions (L7): for each position from , we use the linear approximation to identify promising token substitutions (L8-9), filter irreversible tokenizations (L10), and evaluate the exact objective on the remaining candidate triggers, picking the objective-maximizing token substitution to update (L11).
Overall, GASLITE shows superior speed and efficacy in retrieval attacks compared to the state-of-the-art retrieval attack, Cor.Pois. (Zhong et al., 2023), and outperforming strong discrete optimizers (GCG (Zou et al., 2023), ARCA (Jones et al., 2023)), originally used for LLM jailbreaking, but repurposed here for retrieval attacks (Fig. 2, App. F.1; §6).
5. Experimental Setup
@@
We employ an extensive setup to test the susceptibility of popular, leading retrievers to SEO under varied assumptions and compare GASLITE to adequate baselines (see App. E for more details).
Models. We evaluate diverse embedding-based retrievers (detailed list in Tab. 4 in App. E): MiniLM (Wang et al., 2020); E5 (Wang et al., 2022); Arctic (Merrick et al., 2024); Contriever and Contriever-MS (Izacard et al., 2021), ANCE (Xiong et al., 2020); GTR-T5 (Ni et al., 2022); and MPNet (Song et al., 2020b). We select these models based on performance (per retrieval benchmarks (Muennighoff et al., 2023)), popularity (per HuggingFace’s downloads and open-source usage), diverse architectures (i.e., backbone model, pooling method, and similarity function), extensive usage in prior work, and size (specifically with 110M parameters, the size of BERT-base), as efficiency is a desired property when working with large corpora. We also study attacking LLM-based embeddings (Stella-1.5B (Zhang et al., 2025b)) in §6.5.2.
Datasets. Focusing on retrieval for search, we use the MSMARCO passage retrieval dataset (Bajaj et al., 2016), containing a corpus of 8.8M passages and 0.5M real search queries, which we poison and target in attacks, respectively. For , we sample toxic statements from ToxiGen (Hartvigsen et al., 2022), and for concept-specific content, we use GPT4 (OpenAI, 2024a) to create negative statements. We also validate results on the NQ dataset (Kwiatkowski et al., 2019) (App. G.3), and show that attacks on MSMARCO generalize to other datasets (e.g., NQ, FiQA2018, SciFact and Quora; (Muennighoff et al., 2023))).
Our Attack. To simulate worst-case attacks while ensuring passages remain within benign passage length (App. F.3), we evaluate GASLITE for crafting passages where a malicious prefix is fixed, followed by a trigger of length (i.e., ). We extend GASLITE to a multi-budget attack using -means for query partitioning (§4.1). For additional hyperparameters we set , , and , as elaborated in App. F.3.
Attacks. We consider a naiv̈e baseline, and the two established attacks that populate prior work (Zou et al., 2024; Shafran et al., 2025; Zhong et al., 2023; Chaudhari et al., 2024), for comparison. We set all methods to perform an informative attack (). First, as a control, in info Only we attack with the chosen alone (). Second, following a common SEO attack, we use stuffing (Zuze and Weideman, 2013; Zou et al., 2024; Shafran et al., 2025)—i.e., filling the with sample queries (App. E). Third, we employ the Cor.Pois. method (Zhong et al., 2023), a state-of-the-art HotFlip variant for attacking retrieval, as a representative of HotFlip-based attacks recently proposed against retrievers (Zhong et al., 2023; Zou et al., 2024; Chaudhari et al., 2024). We use the original implementation of Cor.Pois., while allowing the attack to operate under its more permissive threat model where the attacker can access the corpus and to ground truth passages (i.e., the answers to the targeted queries). For fair evaluation, all methods perform query partitioning using -means (§4.1) and share length ().
Metrics. As the informativeness adversarial objective is inherently achieved in the attacks ( serves as a prefix in crafted passages), we measure the attack success in terms of visibility. To this end, we adopt the well-established metric of appeared@k (Zhong et al., 2023; Song et al., 2020a), measuring the proportion of queries for which at least one adversarial passage () appears in the top- results; we typically set =10, per common search apps (e.g., the first page of Google search commonly displays 10 results), taking measurements over held-out queries (except in Knows All attacks in §6.1).
6. Evaluating the Susceptibility of Embedding-based Search
In what follows, we evaluate our attack in three settings (§6.1–6.3), each corresponding to a different type and attacker knowledge of the targeted query distribution (per §3). We chiefly focus on the Knows What setting (§6.2), as it better reflects realistic SEO. Then, we evaluate attack generalizability (across different query sets, and to additional architecture) and transferability (across models) (§6.5) and assess results using the hypothetical perfect attack (§6.4).
6.1. “Knows All” Adversary
@@ 6. Evaluating the Susceptibility of Embedding-based Search
Results. Tab. 1 shows that while the content alone ( Only) never appears in top-10 results, merely appending the query (stuffing) boosts visibility to 30% average appeared@10. Cor.Pois. (Zhong et al., 2023) underperforms, sometimes worse than the naïve stuffing; we find it is mainly, albeit not only, due to generating adversarial tokens that, once decoded into text and tokenized to model input, result in vastly different tokens than the ones optimized (see App. H). Importantly, carefully designing the suffix with GASLITE renders it optimally visible, consistently ranked as the top-1 passage for each query; we attribute this to GASLITE’s passages achieving an exceptionally high vector similarity with the target query (Tab. 13)—for some dot-product models (e.g., Contriever) this similarity is twice that of baselines, a phenomenon we discuss later (§9).
6.2. “Knows What” Adversary
@@
Setup. We test attacks’ ability to increase visibility of specific (e.g., a negative Harry Potter review) across queries on a targeted concept (e.g., Potter). To test this setting, we choose eight recurring concepts of varying semantics and frequency, from MSMARCO (see App. E.4). Per our threat model (§3), we employ a sample (50%) of concept-related queries—albeit this can be relaxed by generating synthetic queries (App. G.6)—leaving a held out set (50%) of concept-related queries for evaluation with varying budget sizes (). Tab. 15 and Tab. 16 list examples of crafted passages.
@@
Results. Fig. 3 shows GASLITE outperforms all baselines, increasing appeared@10 by 40%. Unlike single-query attacks (§6.1), stuffing fails here (except with Contriever; §9) highlighting the increased attack difficulty compared to Knows All. Note that GASLITE remains superior even when evaluated under a more permissive, less realistic threat model, measuring the success directly on the crafted input tokens instead of text (Fig. 19, App. G.2). Further, GASLITE’s superiority is also evident when controlling for the query-partitioning method, by comparing attacks on single passage budget () (Tab. 12 in App. G.2). Fig. 4 demonstrates that attack success increases along the attack budget, with insertion of =10 passages to 8.8M-sized corpus (a poison rate of 0.0001%) sufficing for 50% avg. appeared@10 on 6/9 retrievers for unknown concept-related queries.
@@
Consistent with Zhong et al. (Zhong et al., 2023), we find Contriever models highly susceptible (Figs. 3–4), with a single adversarial passage achieving 100% appeared@10. Other models show varying success (see §9), yet GASLITE’s results approach the optimal solution of Eq. 2 (see §6.4). Last, we observe GASLITE attains 100% appeared@100 on all models for (Fig. 21, App. G.2), indicating substantial content promotion, even if not into the top-10 results.
6.3. “Knows [Almost] Nothing” Adversary
@@
Setup. We test attacks’ potency when targeting general, unknown queries from a wide and diverse query distribution; a challenging setting Zhong et al. (Zhong et al., 2023) studied. Here, we randomly sample 5% of MSMARCO’s training queries (25K queries), made available for attacks with budgets . We then evaluate on MSMARCO’s entire, diverse evaluation queries (7K), held out from the attack.
@@
Results. GASLITE markedly outperforms the baselines (Tab. 17): naïve methods ( only, stuffing) fail to achieve any visibility in top-10 results (or even in top-100) and Cor.Pois. achieves appeared@10 of 6% (Tab. 17). GASLITE’s performance significantly increases along the budget size, from 0% appeared@10 for a single-passage budget (Fig. 22, App. G.3) to appeared@10 of 5%-20% with =100, for most models (Fig. 5). The potential performance for a further increased budget size is explored in §6.4. Consistent with Knows What evaluation (§6.2), we observe variance in attack success across models (see §9).
6.4. The perfect Attack